# Introducing MeMo: A Multimodal Dataset for Memory Modelling in Multiparty Conversations

### v1

@@ Acknowledgements
This material is based upon work supported by Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture, and Science through the Netherlands Organisation for Scientific Research; European Commission funded project “Humane AI: Toward AI Systems That Augment and Empower Humans by Understanding Us, our Society and the World Around Us” (grant # 820437); the National Science Foundation (NWO) under Grant No. (1136993) and Grant No. (024.004.022); TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the supporting institutions. The support is gratefully acknowledged.
@@
The contributions in the work presented in this paper were the following (CRediT taxonomy). Maria Tsfasman: conceptualisation, methodology, resources, data curation, writing - original draft, writing - review & editing, visualisation, project administration, software, formal analysis, investigation. Bernd Dudzik: writing - original draft, writing - review & editing, project administration, supervision. Kristian Fenech: conceptualisation, methodology, software, data curation, funding acquisition, writing - review & editing. Andras Lorincz: conceptualisation, resources, funding acquisition. Catholijn M. Jonker: conceptualisation, methodology, writing - review & editing, funding acquisition. Catharine Oertel: [to be added by Catha].
@@
Maria Tsfasman (a.k.a Masha) is currently pursuing a PhD at TU Delft. Her interdisciplinary memory research aims to enhance machine understanding of humans and use it to gain insights into human cognition. She holds a BSc in Linguistics from HSE University, Moscow, and a MSc in AI with distinction from Radboud University, Nijmegen. She is proud to have been a part of ISIR, Sorbonne University, and IRCN, University of Tokyo throughout her studies. |
Bernd Dudzik received the bachelor’s degree in online media from Furtwangen University, Germany, and the master’s degree in media technology from Leiden University. He has received a doctoral degree with the Intelligent Systems Department, Delft University of Technology. His current research focuses on building computational models of when and how personal memories influence individuals’ emotional experiences of media content. He is currently a member of the Association for the Advancement of Affective Computing. |
@@
Catharine Oertel |

### v2

@@ Acknowledgements
This material is based upon work supported by Delft Institute for Values; Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture, and Science through the Netherlands Organisation for Scientific Research; European Commission funded project “Humane AI: Toward AI Systems That Augment and Empower Humans by Understanding Us, our Society and the World Around Us” (grant # 820437); the National Science Foundation (NWO) under Grant No. (1136993) and Grant No. (024.004.022); TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the supporting institutions. The support is gratefully acknowledged.
@@
The contributions in the work presented in this paper were the following (CRediT taxonomy). Maria Tsfasman: conceptualisation, methodology, resources, data curation, writing - original draft, writing - review & editing, visualisation, project administration, software, formal analysis, investigation. Bernd Dudzik: writing - original draft, writing - review & editing, project administration, supervision. Kristian Fenech: conceptualisation, methodology, software, data curation, funding acquisition, writing - review & editing. Andras Lorincz: conceptualisation, resources, funding acquisition. Catholijn M. Jonker: conceptualisation, methodology, writing - review & editing, funding acquisition. Catharine Oertel: conceptualisation, methodology, resources, data curation, writing - original draft, writing - review & editing, project administration, funding acquisition.
@@
Maria Tsfasman (a.k.a Masha) Her interdisciplinary memory research aims to enhance machine understanding of humans and gain insights into human cognition. She holds a BSc in Linguistics from HSE University, Moscow, and a MSc in AI with distinction from Radboud University, Nijmegen. Before her PhD, she is proud to have been a research assistant at ISIR, Sorbonne University, and IRCN, University of Tokyo. |
Bernd Dudzik is an Assistant Professor in the Department of Intelligent Systems at TU Delft. His research interests focus on Affective Computing and AI-Human collaboration. In particular, his work explores context-sensitive approaches for multimodal modeling of human cognitive-affective processes (e.g., memory recollection or cognitive appraisals) during everyday interactions. Bernd is an active member of the AAAC and Associate Editor for the IEEE Transactions on Affective Computing. |
@@
Catharine Oertel received her PhD degree from the Royal Institute of Technology (KTH), Sweden. After completing her postdoctoral position at EPFL, Switzerland she joined TU Delft as an assistant professor in 2019. Her research focuses on long-term human agent interaction where she particularly focuses on memory modelling. |